# Diversifying the Expert Knowledge for Task-Agnostic Pruning in Sparse Mixture-of-Experts

### v1

@@ 4.3 Empirical Analysis on Expert Hints
4.4 Ablation Study and Discussion
On the number of samples. In our work, we augment the data to improve the accuracy of similarity estimation. While this technique is typically used in scenarios with limited data, a natural question arises:
Would increasing the amount of data help preserve performance after pruning?
@@ 4.4 Ablation Study and Discussion
We conduct experiments with different numbers of samples to answer this question. As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$. It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
@@
On the choice of distance function. We also considered other distance functions, such as cosine similarity, to evaluate expert similarity during the discovery stage. However, it performed worse compared to our mean-squared-error function, showing a drop of $0.9\%$ on MMLU, $3.3\%$ on RTE, and $0.2\%$ on BoolQ.
@@ 5 Conclusion and Future Work
However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating significant potential for further pruning. Future work will aim to push the compression of MoE layers to the limit.
@@ Appendix A Evaluation on the expert similarity
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. A1 and table 2. We summarize the observation as follows,

### v2

@@ 4.5 Empirical Analysis on Expert Hints
4.6 Ablation Study and Discussion
Comparison with the greedy search. To further demonstrate the superiority of our proposed expert pruning strategy, we employ a greedy search approach to identify the optimal candidates in each MoE layer for pruning and compare the results with those obtained using our method. Specifically, we enumerate all possible combinations of removing two experts in each MoE layer, evaluate the pruned models on the MMLU task, and record the best combination for each layer. We then merge these results to prune the model across all layers. The model pruned using the greedy search approach achieves a performance of $62.22\%$, while our data-centric strategy achieves $61.19\%$, showing a comparable result. We provide the full enumeration results in section A.1.
On the number of samples. We conduct experiments to study the effect of the used sample in data-centric pruning methods. As shown in table 3, when we increase the number of samples from $128$ to $512$, we can see a noticeable performance improvement with up to $0.6\%$. It should be noted that when we scale up the number of samples to $2,048$, the average performance dramatically drops to $61.6\%$.
@@ 5 Conclusion and Future Work
However, current methods do not account for the uniqueness of different MoE layers during the pruning process, often setting the same number of experts to be reduced across all layers. Our analysis reveals that some experts have low visitation frequencies, indicating a significant potential for further pruning.
@@ Appendix A Evaluation on the expert similarity
Following the same setting in section 3.2, we conduct experiments on evaluating the expert similarity in all MoE layers of Mixtral-8x7B and Mixtral-8x22B. The results are respectively depicted in fig. A1 and fig. A2. We summarize the observation as follows,
@@
A.1 Enumeration on the expert pruning
We present the full greedy search result on Mixtral-8x7B in fig. A3. In detail, we first enumerate all the possible combinations of dropping $2$ experts layer by layer, and then evaluate the model on the MMLU task. For example, $66.19$ in the first row and third column in layer 0 indicates performance while dropping the first and third expert in layer 0 of the Mixtral-8x7B model.
Although dropping most of the combinations on different layers only leads to a minor performance drop, we can notice that it could cause the model to crash when the fourth expert in layer 1 is involved during the pruning process.
Figure A3: Enumeration on dropping two experts.
A.2 Results on different strategies for pruning
While we report the best performance using our data-centric and model-centric strategies to prune the MoE models, we detail more results on pruning the Mixtral-8x7B and Mixtral-8x22B.
For model centric, we show the results with vectorized weight and surrogate weight strategies to discover the similar experts. We uniformly merge the experts and their routers for model pruning. For data centric, after we discover similar experts, we respectively use the learning strategy to merge the experts with weights (Learn), or only maintain the expert with the maximum visiting frequency (Max).
Table A1: Results on pruning the Mixtral-8x7B and Mixtral-8x22B from 8 experts to 6 and 4 experts in each MoE layer. We present the results of our four strategies, namely 1) Vectorized and surrogate $\theta$: prunable experts discovery using vectorized or surrogate weights and merging with uniform coefficients; 2) Learn: prunable experts discovery using vanilla data and learn coefficients to merge based on eq. 7; 3) Max: maintaining the expert in each discovered group with the maximum visiting frequency. The first and second columns respectively indicate the results on pruned model with $6$ and $4$ experts.
Dataset MMLU BoolQ OpenBookQA RTE Average
Method humanities social science stem Other
Mixtral-8x7B [14] 60.5 77.8 58.9 74.2 85.4 34.4 71.1 66.0
Model-centric Vectorized $\bf{\theta}$ 54.4/48.1 70.2/58.5 51.8/45.2 66.8/55.2 85.6/83.7 31.4/26.2 68.9/62.4 61.3/54.2
[…]
Max 56.4/47.6 71.9/58.6 52.0/42.9 66.9/55.7 85.1/82.8 35.2/28.8 70.4/66.8 62.6/53.7
Mixtral-8x22B [14] 68.6 84.1 67.1 78.7 87.9 0.358 71.2 70.4
Model-centric Vectorized $\bf{\theta}$ 26.7/24.2 28.5/21.7 26.9/21.3 31.7/23.8 62.0/53.6 19.6/11.4 52.0/53.1 35.3/29.9
[…]
Max $\bf{\theta}$ 56.8/47.1 69.7/56.4 50.9/42.2 66.0/55.0 86.9/83.8 32.6/27.0 68.5/64.6 61.6/53.7